out of 5 male worker and 7 female worker of a factory, a task force consisting of 2 male workers and 3 female workers is to be formed.in how many ways can this be done if one particular female must be a member
Answer by Boreal(15235)

If one specific female has to be a member, we now have a group of 4 to be chosen with 5 men and now 6 women. Two of the five have to be men, and two of the 6 women have to be chosen, which will give the desired result.
The number of ways that can occur is 5C2*6C2=10*15=150